OpenStudy (anonymous):

let f(x)=x+5 what is f(x+3)

hartnn (hartnn):

to get f(x+3) just replace 'x' in f(x) by 'x+3'

OpenStudy (anonymous):

x + 8

OpenStudy (anonymous):

f(x) = x + 3 + 5 f(x) = x + 8

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I have another let f(t)=t^2 what is f(t-4)

hartnn (hartnn):

just replace t by t-4 what do u get when you put t-4 instead of t in t^2 ?

OpenStudy (anonymous):

(t-4)^2

hartnn (hartnn):

that would be correct :)